Operations Report – January 15th


Operations Report

Crew 219 Operations Report 15-JAN-2020

Sol: 10

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 16:00

To what time this morning: 06:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 76%

10:00 = 100%; 11:30 = 98%; 12:30 = 95%; 14:00 = 90%; 16:10 = 83%; 18:00 = 78%

Diesel Reading – 60%

Station Propane Reading- 35%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147714,1

Water (static tank) (gal) – 328

Static to loft Pump used –Yes

Water in Green Hab (gal): 161.21

Water in Science Dome (gal): 0

Toilet Tank Emptied –No

Diemos rover used: Still in the Work shop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

Starting Hours: 125.6

Beginning Charge: 100%

Ending Hours: 126.1

Ending Charge: 51%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 131.6

Beginning Charge: 100%

Ending Hours: 132.0

Ending Charge: 51%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used by Outpost for the Post Office and to
acquire an oil pump for the generator.

CrewCar used and why, where? In the shop.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: All suits have been tested for battery
performance. All suit batteries output higher voltage than the
replacement batteries. Suits 1 and 6 have slightly lower air flow on
the right fan, however, this did not cause any major problems during

Summary of Hab: Nominal

Summary of Science Dome operations: Temperature averaging 7°C during
the day and 0°C at night.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support:

Copyright © The Mars Society. All rights reserved. | Main Site